OpenStudy (anonymous):

-5126=-6-5(3x+22)^(5/3) solve

OpenStudy (l):

-5126 = -6 - 5 . (3x+22)^(5/3) -5120 = - 5 . (3x+22)^(5/3) 1024 = (3x+22)^(5/3) 1024^(3/5) = (3x+22)^(5/3)^(3/5) 64 = 3x + 22 42 = 3x -> x = 14 *